OpenStudy (anonymous):

Write a rule for the arithmetic sequence. 5, 56, 107, 158, ...

OpenStudy (mathstudent55):

What is the common difference?

OpenStudy (mathstudent55):

To find the common difference of an arithmetic, just subtract a term from the next term.

OpenStudy (anonymous):

the answers are, hold on.

OpenStudy (anonymous):

A. A(n) = 5 + 51(n - 1) B. A(n) = 5 + 51n C. A(n) = 51 + 5(n - 1) D. A(n) = 51 + 5n

OpenStudy (mathstudent55):

Well, the sequence does start with 5, so the first 2 choices are a possibility. The 3rd and 4th choices can be eliminated.

OpenStudy (mathstudent55):

For the first term, n = 1. Using choice A., you get A(1) = 5 + 51(1 - 1) = 5 + 51(0) = 5 + 0 = 5. The first term is correct. Choice A is still possible. Now let's try the first term with choice B. Using choice A., you get A(1) = 5 + 51(1) = 5 + 51 = 5 + 51 = 56 This is not the first term, so choice B, is wrong. The answer is A.